3. a. Use the Henderson-Hasselbach equation to determine the ratio of HCO3- to H2CO3 in the patient’s blood 10 hours after aspirin ingestion.

b. How does this compare to the ratio of HCO3- to H2CO3 in normal blood (indicate value used from the range given in the table). (pKa1=6.35). c. Can the H2CO3/HCO3- system serve as an effective buffer in this patient? Explain
Answer #1

Q3

assume pH of stomach = 1.5

then

pH = pKa + log(HCO3-/H2CO3)

1.5 = 6.35 + log(HCO3-/H2CO3)

HCO3-/H2CO3 = 10^(1.5-6.35) = 0.000014125

ratio = 0.000014125

b)

compare vs. ratio in blood

assume pH blood = 7.4 approx

pH = pKa + log(HCO3-/H2CO3 )

7.40 = 6.35 + log(HCO3-/H2CO3 )

HCO3-/H2CO3 = 10^(7.40 - 6.35 ) = 11.220

ratio = 11.220

there is much more HCO3- in ionic state in blood than in stomach